A)
1 Find the Laurent series for 22 +22 for 0 < 121 < 2
B)
Find the Laurent series for (z+2)}(3-2) for 2 – 3) > 5
C)
1 Find the Laurent series for z2(z-i) for 1 < 12 – 11 < V2
Answer #1

SWh에 A) let f(2)= 전 <시리<2 저2 02: 뭘시 2 나르 2 그글 자 [ 대를게 <월의 굴 다중 - ② - + - which is the required Laurent series for f(Z). 2근 9i, e glu = [(- 1 4 0 2 ) + (- 2 + 30 - ] g(u) un 7 1 13 or gez) 7 (7-3) which is the required Laurent series for g(2). (2-3)2* H(2)=-6(1+)7 (1+6)-(187)(*** [ Chok. 12 ana INI2V2 1ural = 10,21, (1) V2 21 i b6 + z- -} + - 4 63 + --} & {- {1 - 2 + 3 262